Schedule of Cash Payments for a Service Company SafeMark Financial Inc. was organized on February 28. Projected selling and administrative expenses for each of the first three months of operations are as follows: March $105,300 April 97,900 May 89,100 Depreciation, insurance, and property taxes represent $22,000 of the estimated monthly expenses. The annual insurance premium was paid on February 28, and property taxes for the year will be paid in June. 60% of the remainder of the expenses are expected to be paid in the month in which they are incurred, with the balance to be paid in the following month. Prepare a schedule indicating cash payments for selling and administrative expenses for March, April, and May. SafeMark Financial Inc. Schedule of Cash Payments for Selling and Administrative Expenses For the Three Months Ending May 31 March April May March expenses: Paid in March $ Paid in April $ April expenses: Paid in April Paid in May $ May expenses: Paid in May Total cash payments $ $ $
Answer #1

Answer-

SAFEMARK FINANCIAL INC.
SCHEDULE OF CASH PAYMENTS FOR SELLING AND ADMINISTRATIVE EXPENSES
FOR THE THREE MONTHS ENDING MAY 31
PARTICULARS MARCH APRIL MAY
$ $ $
March expenses:-
Paid in March ($105300-$22000)*60%=$49980
Paid in April ($105300-$22000)*40%=$33320
April expenses:-
Paid in April ($97900-$22000)*60%=$45540
Paid in May ($97900-$22000)*40%=$30360
May expenses:-
Paid in May ($89100-$22000)*60%=$40260
Total cash payments $ 49980 78860 70620